Solve for x: 4cos6x-sin6x=cos^3x+3cos2x

Solve for x:
4cos6x-sin6x=cos^3x+3cos2x

Grade:

1 Answers

Sher Mohammad IIT Delhi
9 years ago
4cos6x-sin6x=cos^3x+3cos2x

for x=0,.
4-0=1+3
means x=0 is a solution

for x=pi
4- 0= -1 +3 , not a solution.

hence only 2npi , is solution of the equation where n=1,2,3..............